OpenStudy (anonymous):

Write the quadratic equation in general form. (x - 3)^2 = 0 is it x^2- 6x + 9 = 0?

OpenStudy (lgbasallote):

\[(x-3)^2 =0\implies (x-3)(x-3) = 0\implies x^2 - 3x - 3x + 9 = 0 \]\[\implies x^2 - 6x + 9\]

OpenStudy (anonymous):

so i'm right ya??:)

OpenStudy (anonymous):

NO. You were wrong...

OpenStudy (lgbasallote):

definitely